OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[\sqrt{72}-\sqrt{147}-5\sqrt{8}+3\sqrt{48}\]

OpenStudy (anonymous):

a.-2sqrt35 b.sqrt5 c.-8+5sqrt6 d.-4sqrt2+5sqrt3

OpenStudy (anonymous):

\[\sqrt{72}=\sqrt{36\times2}=6\sqrt{2}\] \[\sqrt{147}=\sqrt{49\times3}=7\sqrt{3}\] \[\sqrt{8}=\sqrt{4\times2}=2\sqrt{2}\] \[\sqrt{48}=\sqrt{16\times3}=4\sqrt{3}\] complete the answer.

OpenStudy (anonymous):

ok give me onw minute to figure it out

OpenStudy (anonymous):

I got D. -4sqrt2+5sqrt3 is that correct @terenzreignz

terenzreignz (terenzreignz):

Yup, it is :)

OpenStudy (anonymous):

can you check one more for me in a minute please

OpenStudy (anonymous):

\[6\sqrt[3]{5}-\sqrt[4]{5}+2\sqrt[3]{5}+4\sqrt[4]{5}\]

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

a.11sqrt5 b.\[8\sqrt[3]{5}+3\sqrt[4]{5}\] c.11sqrt10 d.\[8\sqrt[3]{10}+3\sqrt[4]{10}\] i got A

OpenStudy (anonymous):

is my answer correct @Jhannybean and @terenzreignz

terenzreignz (terenzreignz):

Unfortunately no :) Why complicate things? Just combine like terms :D

OpenStudy (anonymous):

oh k i will try that

OpenStudy (anonymous):

i am confused icombined like terms and my answer doesnt match any of my choices

OpenStudy (jhannybean):

\begin{align} \large \sqrt{72}-\sqrt{147}-5\sqrt{8}+3\sqrt{48} &= \large 6\sqrt{2} - 7\sqrt{3} -5(2\sqrt{2})+3(4\sqrt{3}) \\ &=\large \color{red}{6\sqrt{2}} -\color{blue}{7\sqrt{3}} -\color{red}{5(2\sqrt{2})} +\color{blue}{3(4\sqrt{3})} \\ \end{align}

OpenStudy (jhannybean):

Match the colors, simplify the highlighted stuff.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i got\[8\sqrt[3]{10}+3\sqrt[4]{10}\]

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

\[\large \color{red}{6\sqrt{2}} -\color{blue}{7\sqrt{3}} -\color{red}{5(2\sqrt{2})} +\color{blue}{3(4\sqrt{3})}\]\[\large (6\sqrt{2}-10\sqrt{2})+(12\sqrt{3}-7\sqrt{3})\]Can you simplify this forme?

OpenStudy (anonymous):

do i distribute?

OpenStudy (jhannybean):

No...just simplify what is inside the parenthesis. You cannot combine the stuff with the \(\sqrt{2}\) attached to them with \(\sqrt{3}\) unless you multiply them together, we're not using that operant here so combining them is not an option. Think of \(\sqrt{2}\) and \(\sqrt{3}\) as x and y values. They cannot be combined here :)

terenzreignz (terenzreignz):

Using colours to emphasise? That's my play, @Jhannybean >:( LOL JK

OpenStudy (jhannybean):

xD

OpenStudy (anonymous):

-4sqrt2+5sqrt3

OpenStudy (jhannybean):

Good job : )

OpenStudy (anonymous):

butthat isnt one of my answer choices

OpenStudy (anonymous):

these are my answer choices; a.11sqrt5 b.85 √ 3 +35 √ 4 c.11sqrt10 d.810 − − √ 3 +310 − −

OpenStudy (anonymous):

d i supose to be \[8\sqrt[3]{10}+3\sqrt[4]{}\]

OpenStudy (anonymous):

and bis supose to be \[8\sqrt[3]{5}+3\sqrt[4]{5}\]

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

That's not even one of your answer choices... a.-2sqrt35 b.sqrt5 c.-8+5sqrt6 d.-4sqrt2+5sqrt3

OpenStudy (anonymous):

the only other one that has a radical in it is b so is that going to be my answer

OpenStudy (jhannybean):

you have radicals in all your solutions??

OpenStudy (anonymous):

your looking at the first question i am doing the second one and i reposted the answer choices

OpenStudy (jhannybean):

Oh,sorry, i didn't see that one.

OpenStudy (jhannybean):

\[\large 6\sqrt[3]{5}-\sqrt[4]{5}+2\sqrt[3]{5}+4\sqrt[4]{5}\] For this one just combine like terms,once again. You got what answer?

OpenStudy (anonymous):

i got d

OpenStudy (anonymous):

i ment b sorry

OpenStudy (anonymous):

\[8\sqrt[3]{5}+3\sqrt[4]{}\]

OpenStudy (jhannybean):

<_<... Lets work it out :P

OpenStudy (jhannybean):

\[\large (6\sqrt[3]{5}+2\sqrt[3]{5})+(4\sqrt[4]{5}-1\sqrt[4]{5})\]

OpenStudy (jhannybean):

And yes, you get \(\large 8\sqrt[3]{5} +3\sqrt[4]{5} \)

OpenStudy (anonymous):

so im correct

OpenStudy (jhannybean):

Yes :)

OpenStudy (anonymous):

thankss